OpenStudy (anonymous):

What is the derivative of y=[x + (x + sin(x)^2)^6]^3 ?

zepdrix (zepdrix):

\[\Large\rm y=\left[x+(x+\sin^2x)^6\right]^3\]Derivative of this?

OpenStudy (xapproachesinfinity):

well use General Power Rule! \((U^n)'=nU^{n-1}U'\) set \(U=x+(x+sin^2x)^6\) \(U'=1+((x+sin^2x)^6)'\) Chain rule again \(6(x+sin^2x)^{5}(1+2cosxsinx)\) then \(U'=1+6(x+sin^2x)^{5}(1+2cosxsinx)\)

OpenStudy (xapproachesinfinity):

Substitute U in the first line

OpenStudy (xapproachesinfinity):

and U' and you are done